# Position: LLMs Can be Good Tutors in Foreign Language Education

### v1

@@ 7 Challenges and Future Directions
LLMs may produce hallucinations (Huang et al., 2023) that can mislead learners and undermine pedagogical goals. This risk intensifies in high-stakes educational environments, where trust and correctness are paramount. Efforts to increase data quality (Long et al., 2024), combine model outputs with structured domain knowledge, and employ rigorous validation mechanisms (including human oversight) are crucial for minimizing such detrimental outcomes.
@@
As LLMs inherit biases from their training data, these systems may produce culturally insensitive or unfair responses, potentially harming students from diverse linguistic and sociocultural backgrounds. Moreover, privacy concerns emerge when collecting and using learner data to personalize instruction. Robust governance frameworks, transparent documentation of data sources, and bias mitigation strategies (Borah & Mihalcea, 2024) can help ensure that LLMs serve as equitable tools in FLE.
@@
LLMs excel at generating language but often lack pedagogical alignment, particularly for tasks requiring developmental sensitivity, motivation, or differentiated instruction. Their general-purpose nature means they do not inherently account for language acquisition theories or curricular standards (Razafinirina et al., 2024). To fully realize their potential, LLM-based tutors should be guided by robust pedagogical frameworks and co-designed with educators who understand learners’ cognitive and developmental needs.
8 Alternative Views
8.1 Task-Specific or Language-Specific Models as Better Alternatives
@@
Some argue that specialized or language-specific models, including classical ML systems with carefully engineered features, can outperform general-purpose LLMs in narrowly defined tasks (e.g., phonetics or grammar drills (Fang et al., 2023)). By focusing on limited objectives, such models avoid the computational overhead and potential inaccuracies of LLMs, which aim to handle a broader range of inputs and contexts (Shen et al., 2024a).
@@
While specialized models may excel in isolated tasks, they lack the flexibility required for comprehensive FLE, which involves cultural nuances, conversations, and evolving learner needs. In contrast, LLMs can be fine-tuned for specific goals while still offering broader linguistic competence (Song et al., 2024a). Additionally, relying on multiple specialized models can be resource-intensive, whereas a well-configured LLM provides a unified framework that balances specialization and scalability.
8.2 Concerns About Over-Reliance on LLMs
Critics warn that over-reliance on LLMs may lead to problems such as generating misleading outputs (Nahar et al., 2024), reducing human interaction, and over-standardizing teaching methods. These issues could undermine the interpersonal and motivational aspects of language learning.
@@ 8.2 Concerns About Over-Reliance on LLMs
These risks highlight the need for balanced integration rather than the replacement of human tutors. LLMs can complement educators by automating repetitive tasks, allowing teachers to focus on individualized support and motivation. Advances in AI safety, such as feedback loops (Tong et al., 2024) and human-in-the-loop systems (Wu et al., 2022), can help minimize inaccuracies (Ho et al., 2024). Additionally, the fine-tuning capabilities of LLMs ensure adaptability, supporting diverse and inclusive learning experiences (Lee et al., 2024b).
9 Conclusion
This position paper emphasizes the transformative potential of LLMs in FLE, positioning them as valuable tutors to complement traditional teaching methods. Through their roles as data enhancers, task predictors, and agents, LLMs can provide adaptive learning experiences across the core skills of listening, speaking, reading, and writing. This paper encourages interdisciplinary exploration to FLE.
Impact Statement
This paper positions LLMs as promising tutors in FLE. By leveraging LLMs, we aim to enhance inclusivity, personalization, and overall accessibility. However, potential risks—such as bias, privacy issues, and over-reliance on AI tutoring—necessitate careful oversight, which have been discussed in the paper. In particular, ensuring equitable data representation, model usage transparency, and human agency preservation remains critical. While careful integration with human expertise is essential, the potential for LLMs to revolutionize FLE is significant, paving the way for more inclusive, efficient, and engaging educational practices. We encourage continuing dialogue and interdisciplinary collaboration to responsibly integrate LLMs into educational ecosystems and mitigate potential harms.
Appendix A Literature Review
We provide an overview of LLM-centric research of FLE presented in Figure 4.
LLMs for Foreign Language Education
LLM as Data Enhancer (Section 4) Data Creation PFQS (Li & Zhang, 2024), MCQG-SRefine (Yao et al., 2024), (Lee et al., 2024c), BF-TC (Liu et al., 2024d), MathVC (Yue et al., 2024), EduAgent (Xu et al., 2024), Generative Students (Lu & Wang, 2024) Data Reformation Book2Dial (Wang et al., 2024b), Slide2Lecture (Zhang-Li et al., 2024), WikiDomains (Asthana et al., 2024), (Freyer et al., 2024), (Day et al., 2025), Anthropological Prompting (AlKhamissi et al., 2024), (Liu et al., 2024a) Data Annotation EXCGEC (Ye et al., 2024), EDEN (Siyan et al., 2024), MEGAnno+ (Kim et al., 2024), PDFChatAnnotator (Tang et al., 2024), Coannotating (Li et al., 2023) LLM as Task Predictor (Section 5) Discriminative TOREE (Zhuang et al., 2024), LLM-KT (Zhan et al., 2024), CLST (Jung et al., 2024), Diallogue-KT (Scarlatos & Lan, 2024), (Neshaei et al., 2024), (Mizumoto & Eguchi, 2023), (Seßler et al., 2024), (Syamkumar et al., 2024) Generative EXPECT (Fei et al., 2023), GEE (Song et al., 2024b), EXCGEC (Ye et al., 2024) FELT (Borges et al., 2024), LLM-as-a-tutor (Han et al., 2024), (Stamper et al., 2024) SocraticLM (Liu et al., 2024b), (Favero et al., 2024), (Xiao et al., 2024) Mixed FABRIC (Han et al., 2023), (Stahl et al., 2024b), ReaLMistake (Kamoi et al., 2024), (Lu et al., 2024) LLM-empowered Agent (Section 6) Abilities FOKE (Hu & Wang, 2024), KnowEdit (Zhang et al., 2024a), (Zhang et al., 2024a), PedCoT (Jiang et al., 2024b), LHP (Sonkar et al., 2024), (Razafinirina et al., 2024), LessonPlanner (Fan et al., 2024), MemoryBank (Zhong et al., 2024), (Hu et al., 2024), vNMF (Jiang et al., 2024a), ChatTutor (Chen et al., 2024), (Zheng et al., 2024) Applications SimClass (Zhang et al., 2024b), MathVC (Yue et al., 2024), BIPED (Kwon et al., 2024), LLM-as-a-tutor (Han et al., 2024), KORLINGS (Lee et al., 2024a), (Liu et al., 2024d)
Figure 4: An overview of LLM-centric research of FLE.
Appendix B Four Phases of Research Roadmap
Stage 1: Rule-based Models (1960s–1990s).
Early solutions relied on handcrafted linguistic rules to process language in tightly constrained scenarios (Grosan et al., 2011; C Angelides & Garcia, 1993). Classical platforms like PLATO (Hart, 1981) and Systran (Toma, 1977) operated effectively for highly structured tasks (e.g., grammar drills) but struggled with complex, context-dependent interactions.
Stage 2: Statistical Models (1990s–2010s).
With the increased availability of digitized corpora, methods such as the early version of Google Translate (Och, 2006) and Dragon NaturallySpeaking (Blair, 1997) pioneered statistical pattern mining. These approaches leveraged large datasets to infer linguistic rules probabilistically, improving scalability yet still lacking deeper semantic understanding.
Stage 3: Neural Models (2010s–2020s).
The advent of deep learning architectures (e.g., RNNs (Yu et al., 2019) and Transformers (Vaswani, 2017)) enabled more robust context modeling, sparking transformative applications like Grammarly (Fitria, 2021) and Duolingo (Vesselinov & Grego, 2012). These systems offered enhanced personalization and feedback, significantly augmenting learners’ writing and reading comprehension.
Stage 4: Large Language Models (2020s–Present).
Today’s LLMs (e.g., ChatGPT (Achiam et al., 2023)) combine massive pre-training corpora with generative capabilities, achieving impressive results in multi-turn dialogue, individualized scaffolding, and multimodal integration. Tools such as Khanmigo (Anand, 2023) demonstrate LLMs’ potential for real-time conversational practice, dynamic content creation, and inclusive educational support at scale.

### v2

@@ 7 Challenges
LLMs may produce hallucinations Huang et al. (2023) that can mislead learners and undermine pedagogical goals. This risk intensifies in high-stakes educational environments, where trust and correctness are paramount. Future directions include enhancing data quality and diversity for training Long et al. (2024), developing techniques to integrate LLM outputs with structured domain knowledge and pedagogical rules, and employing rigorous automated and human-in-the-loop validation mechanisms to minimize such detrimental outcomes and improve the factual grounding of LLM-generated educational content.
@@
As LLMs inherit biases from their training data, these systems may produce culturally insensitive, stereotypical, or unfair responses, potentially harming students from diverse linguistic and sociocultural backgrounds. Moreover, significant privacy concerns emerge when collecting and using learner data to personalize instruction, particularly for K-12 students. Future research must focus on developing robust governance frameworks, transparent documentation of data sources and model behaviors, and advanced bias detection and mitigation strategies Borah and Mihalcea (2024); He and Li (2024) to ensure that LLM-based tools for English Education are equitable, fair, and uphold stringent data protection standards.
@@
LLMs excel at generating fluent language but often lack deep pedagogical alignment, particularly for tasks requiring developmental sensitivity, learner motivation strategies, or differentiated instruction tailored to individual learning needs. Their general-purpose nature means they do not inherently account for established language acquisition theories or specific curricular standards Razafinirina et al. (2024). A crucial future direction is the development of methodologies to better imbue LLMs with pedagogical intelligence. This includes co-designing LLM applications with educators, fine-tuning models on high-quality pedagogical interaction data, and creating architectures that can dynamically adapt to learners’ cognitive states and developmental needs in English language learning.
8 Conclusion
This paper emphasizes the transformative potential of LLMs in English Education, positioning them as valuable tutors to complement traditional teaching methods. Through their roles as data enhancers, task predictors, and agents, LLMs can provide adaptive learning experiences across the core skills of listening, speaking, reading, and writing. This paper encourages continuing dialogue and interdisciplinary collaboration to responsibly integrate LLMs into educational ecosystems.
Limitations
Emphasis on potential over practical implementation barriers.
This paper primarily focuses on the potential of LLMs to serve as effective tutors in English Education, outlining beneficial roles as data enhancers, task predictors, and agents. While we acknowledge the existence of challenges (to be discussed in Appendix 7), a limitation of this position is that the main arguments may not fully capture the considerable practical, socio-economic, and infrastructural hurdles that could impede the equitable and effective implementation of these LLM roles across diverse global educational contexts and resource settings.
Generalizability and contextual adaptation of proposed roles.
We propose three broad roles for LLMs in English Education. However, this paper does not provide an exhaustive analysis of how the efficacy and suitability of LLMs in these roles might vary significantly across different target languages (especially low-resource languages), specific learner demographics (e.g., preschoolers vs. K-12 vs. adult learners, learners with disabilities), diverse cultural contexts, or varying pedagogical philosophies. The general framework presented may require substantial adaptation and further research to be effectively applied in specific English Education scenarios.
Nuances of human-LLM pedagogical interaction.
While advocating for LLMs as tutors that can complement human expertise, this position paper does not delve deeply into the complex dynamics of the pedagogical interactions between learners, LLM-based tutors, human educators, and parents. Critical aspects such as optimizing the collaborative model, designing effective training for educators to leverage LLMs, mitigating risks of learner over-reliance, and ensuring that LLM interactions foster deep learning rather than superficial engagement are multifaceted issues that warrant more extensive investigation than afforded by the scope of this paper.
Acknowledgements
This research is supported in part by NSF under grants III-2106758, and POSE-2346158. This research is also supported by National Natural Science Foundation of China (Grant No.62276154); Research Center for Computer Network (Shenzhen) Ministry of Education, the Natural Science Foundation of Guangdong Province (Grant No.2023A1515012914 and 440300241033100801770); Basic Research Fund of Shenzhen City (Grant No.JCYJ20210324120012033, JCYJ20240813112009013 and GJHZ20240218113603006); The Major Key Project of PCL for Experiments and Applications (PCL2024A08).
Appendix A Literature Review
We provide an overview of LLM-centric research of English Education presented in Figure 4.
{forest}
for tree= forked edges, grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=middle, font=, rectangle, line width=0.8pt, draw=output-black, rounded corners,align=left, minimum width=2em, s sep=6pt, l sep=8pt, , where level=1text width=0.2, where level=2text width=0.2font=, where level=3font=, where level=4font=, where level=5font=, [LLMs for English Education, middle,rotate=90,anchor=north,edge=output-black [LLM as Data Enhancer
(Section 4),middle,anchor=west,edge=output-black, text width=0.20[Data Creation, middle, text width=0.14edge=output-black [ PFQS Li and Zhang (2024), MCQG-SRefine Yao et al. (2024),
Lee et al. (2024c), BF-TC Liu et al. (2024d), MathVC Yue et al. (2024),
EduAgent Xu et al. (2024), Generative Students Lu and Wang (2024) , leaf, text width=0.48edge=output-black] ] [Data Reformation, middle, text width=0.14edge=output-black [ Book2Dial Wang et al. (2024b), Slide2Lecture Zhang-Li et al. (2024),
WikiDomains Asthana et al. (2024), Freyer et al. (2024),
Day et al. (2025), Anthropological Prompting AlKhamissi et al. (2024),
Liu et al. (2024a) , leaf, text width=0.48edge=output-black] ] [Data Annotation, middle, text width=0.14edge=output-black [ MEGAnno+ Kim et al. (2024), EDEN Siyan et al. (2024),
PDFChatAnnotator Tang et al. (2024), Coannotating Li et al. (2023) , leaf, text width=0.48edge=output-black] ] ] [LLM as Task Predictor
(Section 5),middle,anchor=west,edge=output-black, text width=0.20[Discriminative, middle, text width=0.14edge=output-black [ TOREE Zhuang et al. (2024), LLM-KT Zhan et al. (2024),
CLST Jung et al. (2024), Diallogue-KT Scarlatos and Lan (2024),
Neshaei et al. (2024), Mizumoto and Eguchi (2023), Song et al. (2025),
Seßler et al. (2024), Syamkumar et al. (2024) , leaf, text width=0.48edge=output-black] ] [Generative, middle, text width=0.14edge=output-black [ RECIPE Han et al. (2023a), EXPECT Fei et al. (2023),
GEE Song et al. (2024b), EXCGEC Ye et al. (2024),
FELT Borges et al. (2024), LLM-as-a-tutor Han et al. (2024),
SocraticLM Liu et al. (2024b), Stamper et al. (2024),
Favero et al. (2024), Xiao et al. (2024), Scarlatos et al. (2025) , leaf, text width=0.48edge=output-black] ] [Mixed, middle, text width=0.14edge=output-black [ FABRIC Han et al. (2023b), eRevise+RF Liu et al. (2025),
ReaLMistake Kamoi et al. (2024), Stahl et al. (2024b), Lu et al. (2024) , leaf, text width=0.48edge=output-black] ] ] [LLM-empowered Agent
(Section 6),middle,anchor=west,edge=output-black, text width=0.20[Abilities, middle, text width=0.14edge=output-black [ FOKE Hu and Wang (2024), KnowEdit Zhang et al. (2024a),
PedCoT Jiang et al. (2024b), LessonPlanner Fan et al. (2024),
MemoryBank Zhong et al. (2024), vNMF Jiang et al. (2024a),
ChatTutor Chen et al. (2024), LHP Sonkar et al. (2024),
Razafinirina et al. (2024), Zhang et al. (2024a),
Hu et al. (2024), Zheng et al. (2024), Shi et al. (2025) , leaf, text width=0.48edge=output-black] ] [Applications, middle, text width=0.14edge=output-black [ SimClass Zhang et al. (2024b), MathVC Yue et al. (2024),
BIPED Kwon et al. (2024), LLM-as-a-tutor Han et al. (2024),
KORLINGS Lee et al. (2024a), Liu et al. (2024d) , leaf, text width=0.48edge=output-black] ] ] ]
Figure 4: An overview of LLM-centric research of FLE.
Appendix B Future Directions
Establishing Robust Evaluation Frameworks.
A significant challenge in leveraging LLMs for English Education is the current lack of widely accepted and easily implementable evaluation frameworks to assess the quality of LLM-based teaching interactions and outcomes. Existing metrics often focus on linguistic correctness or task completion Tan et al. (2024); Macina et al. (2025) rather than pedagogical efficacy or impact on learning Chiang et al. (2024). Future work should prioritize the development of standardized evaluation methodologies, including comprehensive benchmarks and nuanced metrics that capture both the accuracy of linguistic information and the pedagogical value of LLM interventions. This will be essential for comparing different systems and guiding iterative improvements.
Integrating with Modern Standardized Educational Frameworks.
English language learning is often governed by established standards and frameworks, such as the Common European Framework of Reference for Languages (CEFR) [^1] https://www.coe.int/en/web/portal/home or Common Core State Standards (CCSS) [^2] https://corestandards.org/. For LLM-based tools to be truly effective and gain acceptance, their outputs and interaction patterns should align with these existing frameworks. Future technical development should focus on enabling LLMs to reference, interpret, and operate consistently within these standards Nicholls et al. (2024); Imperial et al. (2024). This includes generating proficiency-level-appropriate content, providing feedback that corresponds to specific framework descriptors, and assisting learners in achieving standardized learning objectives, thereby enhancing usability, conformity, and trustworthiness among educators and learners.
Fostering Human-AI Collaboration in Pedagogy.
While LLMs offer transformative potential, it is unlikely they will completely replace human teachers in English Education in the foreseeable future. Instead, the most promising path involves developing sophisticated human-AI collaborative educational technologies KIM et al. (2025). Future research should explore how LLMs can best function as assistive tools that augment, rather than supplant, the capabilities of human educators Shojaei et al. (2025). This includes designing intuitive interfaces for teachers to guide, customize, and oversee LLM-driven activities, investigating teachers’ perspectives on integrating LLMs into their practice, and defining technical benchmarks for when an LLM possesses sufficient acquired skills to reliably assist teachers. The focus must be on a synergistic model where LLMs handle scalable tasks while human teachers provide the crucial elements of empathy, nuanced understanding, and holistic student development.
Appendix C Alternative Views
While this paper supports the use of LLMs in English Education, it is essential to consider alternative perspectives. Below, we discuss two key opposing views and provide counterarguments.
C.1 Task-Specific or Language-Specific Models as Better Alternatives
@@
Some argue that specialized or language-specific models, including classical ML systems with carefully engineered features, can outperform general-purpose LLMs in narrowly defined tasks (e.g., phonetics or grammar drills Fang et al. (2023)). By focusing on limited objectives, such models avoid the computational overhead and potential inaccuracies of LLMs, which aim to handle a broader range of inputs and contexts Shen et al. (2024a).
@@
While specialized models may excel in isolated tasks, they lack the flexibility required for comprehensive English Education, which involves cultural nuances, conversations, and evolving learner needs. In contrast, LLMs can be fine-tuned for specific goals while still offering broader linguistic competence Song et al. (2024a). Additionally, relying on multiple specialized models can be resource-intensive, whereas a well-configured LLM provides a unified framework that balances specialization and scalability.
C.2 Concerns About Over-Reliance on LLMs
Critics warn that over-reliance on LLMs may lead to problems such as generating misleading outputs Nahar et al. (2024), reducing human interaction, and over-standardizing teaching methods. These issues could undermine the interpersonal and motivational aspects of language learning.
@@ C.2 Concerns About Over-Reliance on LLMs
These risks highlight the need for balanced integration rather than the replacement of human tutors. LLMs can complement educators by automating repetitive tasks, allowing teachers to focus on individualized support and motivation. Advances in AI safety, such as feedback loops Tong et al. (2024) and human-in-the-loop systems Wu et al. (2022), can help minimize inaccuracies Ho et al. (2024). Additionally, the fine-tuning capabilities of LLMs ensure adaptability, supporting diverse and inclusive learning experiences Lee et al. (2024b).